.) The cost structure of a manufacturer of microships is described in the table shown below.

The firm's fixed costs equal $15,000 per day. Calculate the average variable cost, average fixed cost, and average total cost at each output level. (Your answer should be rounded to the nearest cent.)

Output

(microchips per day)

Total Cost of Output

Average Variable Cost ($)

Average Fixed Cost

($)

Average Total Cost

($)

0

$15,000

-

-

-

20,000

65,000

45,000

100,000

70,000

155,000

95,000

225,000

120,000

330,000

145,000

470,000
Answer #1

OUTPUT TOTAL COST AVC AFC ATC 15,000 65,000 100,000 155,000 225,000 330,000 470,000 2.5 1.88 20,000 45,000 70,000 95,000 120,

The average total cost is found by dividing the total cost by the output produced, the average fixed cost also found out by the same manner. It is calculated by dividing the total fixed cost by the output produced. The last one is the average variable cost, it is calculated by dividing the total variable cost by the output produced but here the total variable cost is not directly given and the total variable cost is calculated by deducting the total fixed cost from the total cost.

AVC=TVC/Q

TVC=TC-TFC

ATC =TC/Q

AFC = TFC/Q